# The Open Journal of Astrophysics

### v1

@@ 2.3.1 Projected correlation function $w_{g+}$
The most widely used statistic to measure the projected shape-position correlation functions of IA is $w_{g+}$. We can obtain $w_{g+}$ by integrating $\xi_{g+}(r_{p},\Pi)$ (Eq. (4)) over the line of sight $\Pi$:
@@
where $\Pi_{max}$ is the chosen integration limit, set to half the boxsize in this work, equating to $102.5\mathrm{Mpc}/h$ for TNG300. As the quadrupole (see below) has a higher signal-to-noise (Singh et al., 2024) than $w_{g+}$, we have chosen to show only the quadrupole in Section 3, but some of the $w_{g+}$ measurements are added to Appendix B for comparison to other works. Any quadrupole measurements in this work are also available as $w_{g+}$ measurements upon request.
@@ 2.3.2 Multipoles
$\tilde{\xi}_{jk}^{\ell,s_{ab}}(r)$ $=\frac{2\ell+1}{2}\frac{(\ell-s_{ab})!}{(\ell+s_{ab})!}\int\text{d}\mu_{r}L^{\ell,s_{ab}}(\mu_{r})\xi_{jk}(r,\mu_{r})$ (8)
where $j$ and $k$ can be $g$ or ++; our focus in this work lies with $j=g$ and $k=+$. The previously defined $\xi_{g+}(r,\mu_{r})$ (Eq. (4)) is inserted into the integrand, along with the appropriate associated Legendre polynomials, where $s_{ab}=\ell=2$ are needed for $\xi_{g+}$.
@@
For the measurements of both $w_{g+}$ and $\tilde{\xi}_{g+,2}$ the code MeasureIA  [^1] https://github.com/MarloesvL/measure_IA was developed and used. MeasureIA is publicly available and has been validated against halotools  [^2] https://github.com/astropy/halotools (Hearin et al., 2017) and TreeCorr  [^3] https://github.com/rmjarvis/TreeCorr (Jarvis et al., 2004) for $w_{g+}$ and $w_{gg}$ measurements.
@@ 2.3.3 Covariance
The covariance is estimated using the jackknife method. In this method, the simulation volume is divided into sub-volumes of equal size, and the correlation functions are measured for the entire volume, emitting one sub-volume at a time. This process is sensitive to the number of sub-volumes, or jackknife regions, $N_{\mathrm{jk}}$, chosen, where Hirata et al. (2004) found that $N_{\mathrm{jk}}$ needs to be larger than $N_{\mathrm{bin}}^{2/3}$, where $N_{\mathrm{bin}}$ is the number of radial bins, and that the sub-box-length needs to be larger than the largest scale that is probed. In this study, we are using 10 radial bins and a maximum scale of $20\mathrm{Mpc}/h$, which makes $N_{\mathrm{jk}}=64$ compatible with the above criterion.
@@
$C_{ij}$ $=\frac{N_{\mathrm{jk}}-1}{N_{\mathrm{jk}}}\sum_{n=1}^{N_{\mathrm{jk}}}(\psi^{n}_{i}-\bar{\psi_{i}})(\psi^{n}_{j}-\bar{\psi_{j}})$ (10)
$\mathrm{with}\ \bar{\psi_{i}}=\frac{1}{N_{\mathrm{jk}}}\sum_{n=1}^{N_{\mathrm{jk}}}\psi_{i}^{n}$ (11)
@@
where $i,j$ run over all radial bins and $n$ over all jackknife realisations from $1$ to $N_{\mathrm{jk}}$.
@@ 2.4 Selections
The position sample is defined by all galaxies with $M_{\star}>10^{8.35}\mathrm{M_{\odot}}/h$, which corresponds to $50$ stellar particles in the lowest resolution simulation (TNG300), and can therefore be considered the largest set of resolved galaxies with a consistent definition between simulations. Due to boxsize differences, this selection corresponds to $407180$ galaxies in TNG300, $115208$ in Horizon-AGN and $22279$ in EAGLE, at $z=0$. For the shape sample, we opt for a consistent stellar mass cut of $M_{\star}>10^{9.15}\mathrm{M_{\odot}}/h$, which corresponds to $300$ stellar particles in TNG300, and is therefore the largest consistent set of galaxies with converged shapes (Chisari et al., 2015).
For consistency, many variables are remeasured from the particle data (in TNG300 and EAGLE) using a code developed for this reason, even if similar variables are available in simulation sub-halo catalogues.
In subsequent sections, we use $\kappa_{\mathrm{rot}}$, $|v/\sigma|$, bulge-to-total ratio ($\mathrm{BTR}$) and rir-i colour to split the shape sample into disks and ellipticals (see Section 3.1). $\kappa_{\mathrm{rot}}$ is measured in TNG300 and EAGLE and is defined as follows:
$\kappa_{rot,j}=\frac{\sum^{N_{j}}_{i=1}\left[\frac{1}{2}m_{i}(\hat{\mathbf{L}_{j}}\times\hat{\mathbf{r}_{i}})\cdot\mathbf{v}_{i}\right]}{\sum^{N_{j}}_{i=1}\left[\frac{1}{2}m_{i}\mathbf{v}_{i}\cdot\mathbf{v}_{i}\right]}$ (13)
@@
where the $^\hat{}$ -symbol denotes the direction of the vector and $\bf{v}_{i}$ is the velocity of each particle $i$ in galaxy $j$ relative to the centre of velocity of the galaxy, defined by:
$\mathbf{v}_{i}$ $=\mathbf{\dot{x}}_{i}-\mathbf{\dot{x}}_{j}$ (14)
$\mathbf{\dot{x}}_{j}$ $=\frac{1}{M_{j}}\Sigma^{N_{j}}_{i=1}m_{i}\mathbf{\dot{x}}_{i}$ (15)
@@
with stellar particle velocity, $\mathbf{\dot{x}}_{i}$ and mass-weighted velocity of the stellar particles for each galaxy, $\mathbf{\dot{x}}_{j}$. The angular momentum of the galaxy is given by:
$\mathbf{L}_{j}$ $=\Sigma^{N_{j}}_{i=1}m_{i}\mathbf{r}_{i}\times\mathbf{v}_{i}\ .$ (17)
@@
$|v/\sigma|$ is the rotational velocity divided by the velocity dispersion of a galaxy, measured following the procedure in Dubois et al. (2016). First, we move to a cylindrical coördinate system with the angular momentum (Eq. (17)) as $z$-axis and all coördinates ($r,\theta,z$) relative to the galaxy centre of mass and velocity ($\mathbf{r}_{i}$, $\mathbf{v}_{i}$). Second, the mass-weighted mean of the $\theta$ component of the velocity need to be computed,
$\overline{v_{\theta,j}}=\frac{1}{M_{j}}\Sigma^{N_{j}}_{i=1}m_{i}v_{\theta,i}\ .$ (18)
@@
$\overline{v_{k,j}}$ $=\frac{1}{M_{j}}\Sigma^{N_{j}}_{i=1}m_{i}v_{k,i}$ (19)
$\overline{\Delta v_{k,j}}$ $=\frac{1}{M_{j}}\Sigma^{N_{j}}_{i=1}m_{i}(v_{k,i}-\overline{v_{k,j}})^{2}$ (20)
$\sigma$ $=\sqrt{\frac{\overline{\Delta v_{r,j}}+\overline{\Delta v_{\theta,j}}+\overline{\Delta v_{z,j}}}{3}}$ (21)
where $k$ is $r,\theta$ or $z$ and $i$ runs over the particles of galaxy $j$ as before. $|v/\sigma|$ is then defined as the absolute value of $\overline{v_{\theta,j}}/\sigma$.
The photometric colours used in this work are all from the available catalogues described in Nelson et al. (2017) for TNG, Trayford et al. (2015) for EAGLE and Kaviraj et al. (2017) for Horizon-AGN, and measured to be SDSS-like, mimicking the colour filters of the Sloan Digital Sky Survey (SDSS), including dust extinction.
Figure 1: Normalised distributions of $\mathrm{BTR}$ (top left), $|v/\sigma|$ (top right), $\kappa_{\mathrm{rot}}$ (bottom left) and rir-i (bottom right), which are used to split galaxies according to morphology, in TNG100 (light blue), TNG300 (dark blue), EAGLE (orange) and Horizon-AGN (green) at $z\approx 0$. The variables show a range of distributions with varying features across simulations.
The bulge-to-total ratio (BTR), which is not available for EAGLE and Horizon-AGN, is taken directly from the TNG300 catalogue (Genel et al., 2015), and therefore available for a limited number of galaxies, i.e. $M_{\star}>10^{9.3}\mathrm{M_{\odot}}/h$ at $z=0$ and $M_{\star}>10^{9.8}\mathrm{M_{\odot}}/h$ at higher redshifts. Therefore, $\mathrm{BTR}$ is included only in Section 3.1, and then examined separately in Appendix D. The $\mathrm{BTR}$ defined as 11- the fractional mass of stars in a galaxy with $\epsilon>0.7$, and measured as follows. First, the angular momentum of the stars within 10 times the stellar half-mass radius is computed. Second, the system’s z-axis is aligned with this angular momentum. Then, the stellar particles taken into account are sorted by their binding energy. For every stellar particle, the specific angular momentum, $J_{z}$ is calculated. $J(E)$ is noted, the maximum angular momentum of the stellar particles between 50 before and 50 after the particle in question (in the sorted list). Then, $\epsilon$ is computed for each particle by $\epsilon=J_{z}/J(E)$. Finally, the bulge to total ratio is computed as defined above.
@@ 3.1 Distributions
Figure 1 shows a comparison between the normalised distributions of the bulge-to-total ratio, $\mathrm{BTR}$ (top left panel), $|v/\sigma|$ (top right panel), $\kappa_{\mathrm{rot}}$ (bottom left panel) and the colour rir-i (bottom right panel), which were described in Section 2.4, for the simulations TNG100 (light blue), TNG300 (dark blue), EAGLE (orange) and Horizon-AGN (green) at $z=0$ ($z=0.0556$ for Horizon-AGN), when available. These variables will be used in following sections to split the galaxy samples in disks and ellipticals.
Figure 1 shows that the simulations agree on the ranges of the variable values, but not on how the variables are distributed. For TNG100 and TNG300, these differences can be ascribed to resolution differences that are explored in Appendix A, together with a discussion on the influence of boxsize and resolution on the results.
@@
The differences between TNG, EAGLE and Horizon-AGN cannot be as easily explained as the differences between simulations in the same suite. Apart from varying resolutions and boxsizes, these simulations have different models and subgrid-physics recipies, as described in Section 2.1. Furthermore, where the kinematic parameters, $\kappa_{\mathrm{rot}}$ and $|v/\sigma|$, may be similar between TNG and EAGLE, the distributions of the colours have diverging features, of which only rir-i is shown in Figure 1. TNG displays a bi-modality in its colours, as found in observational surveys, e.g. GAMA (Taylor et al., 2014) and SDSS (Blanton et al., 2003), whereas this feature is not present in EAGLE or Horizon-AGN. Colour bi-modality has been found in Horizon-AGN (Kaviraj et al., 2017), albeit weaker than in observations. As there are many low-mass red galaxies in Horizon-AGN (Kaviraj et al., 2017), our mass cut is a likely cause of the uni-modality seen in Figure 1. EAGLE also contains too many low-mass red galaxies due to limited numerical resolution (Trayford et al., 2015). For a detailed comparison between GAMA and EAGLE, see Trayford et al. (2015).

### v2

@@ 2.3.1 Projected correlation function $w_{g+}$
The most widely used statistic to measure the projected shape-position correlation functions of IA is $w_{g+}$. We can obtain $w_{g+}$ by integrating $\xi_{g+}(r_{p},\Pi)$ (Eq. 4) over the line of sight $\Pi$:
@@
where $\Pi_{max}$ is the chosen integration limit, set to half the boxsize in this work, equating to $102.5\mathrm{Mpc}/h$ for TNG300. As the quadrupole (see below) has a higher signal-to-noise than $w_{g+}$ (Singh et al., 2024), we have chosen to show only the quadrupole in Section 3, but some of the $w_{g+}$ measurements are added to Appendix B for comparison to other works. Any quadrupole measurements in this work are also available as $w_{g+}$ measurements upon request.
@@ 2.3.2 Multipoles
$\tilde{\xi}_{kl}^{\ell,s_{ab}}(r)$ $=\frac{2\ell+1}{2}\frac{(\ell-s_{ab})!}{(\ell+s_{ab})!}\int\text{d}\mu_{r}L^{\ell,s_{ab}}(\mu_{r})\xi_{jk}(r,\mu_{r})$ (8)
where $k$ and $l$ can be $g$ or ++; our focus in this work lies with $k=g$ and $l=+$. The previously defined $\xi_{g+}(r,\mu_{r})$ Eq. 4 is inserted into the integrand, along with the appropriate associated Legendre polynomials, where $s_{ab}=\ell=2$ are needed for $\xi_{g+}$.
@@
For the measurements of both $w_{g+}$ and $\tilde{\xi}_{g+,2}$ the code MeasureIA  [^1] https://github.com/MarloesvL/measure_IA222doi:10.5281/zenodo.17252215 was developed and used. MeasureIA is publicly available and has been validated against halotools  [^3] https://github.com/astropy/halotools (Hearin et al., 2017) and TreeCorr  [^4] https://github.com/rmjarvis/TreeCorr (Jarvis et al., 2004) for $w_{g+}$ and $w_{gg}$ measurements, including covariance.
@@ 2.3.3 Covariance
The covariance is estimated using the jackknife method. In this method, the simulation volume is divided into sub-volumes of equal size, and the correlation functions are measured for the entire volume, omitting one sub-volume at a time. This process is sensitive to the number of sub-volumes, or jackknife regions, $N_{\mathrm{sub}}$, chosen, where Hirata et al. (2004) found that $N_{\mathrm{sub}}$ needs to be larger than $N_{\mathrm{bin}}^{3/2}$, where $N_{\mathrm{bin}}$ is the number of radial bins, and that the sub-box-length needs to be larger than the largest scale that is probed. In this study, we are using 10 radial bins and a maximum scale of $20\mathrm{Mpc}/h$, which makes $N_{\mathrm{sub}}=64$ compatible with the above criterion.
@@
$C_{kl}$ $=\frac{N_{\mathrm{sub}}-1}{N_{\mathrm{sub}}}\sum_{n=1}^{N_{\mathrm{sub}}}(\psi^{n}_{k}-\bar{\psi_{k}})(\psi^{n}_{l}-\bar{\psi_{l}})$ (10)
$\mathrm{with}\ \bar{\psi_{k}}=\frac{1}{N_{\mathrm{sub}}}\sum_{n=1}^{N_{\mathrm{sub}}}\psi_{k}^{n}$ (11)
@@
where $k,l$ run over all radial bins and $n$ over all jackknife realisations from $1$ to $N_{\mathrm{sub}}$.
@@ 2.4 Selections
The position sample is defined by all galaxies with $M_{\star}>10^{8.35}\mathrm{M_{\odot}}/h$, which corresponds to $50$ stellar particles in the lowest resolution simulation (TNG300), and can therefore be considered the largest set of resolved galaxies with a consistent definition between simulations. Due to boxsize differences, this selection corresponds to $407180$ galaxies in TNG300, $115208$ in Horizon-AGN and $22279$ in EAGLE, at $z=0$ (see Table 2 for all sample sizes). For the shape sample, we opt for a consistent stellar mass cut of $M_{\star}>10^{9.15}\mathrm{M_{\odot}}/h$, which corresponds to $300$ stellar particles in TNG300, and is therefore the largest consistent set of galaxies with converged shapes (Chisari et al., 2015). In subsequent sections, we use $\kappa_{\mathrm{rot}}$, $|v/\sigma|$, bulge-to-total ratio ($\mathrm{BTR}$) and rir-i colour to split the shape sample into disks and ellipticals (see Section 3.1). The motivation for using these variables is based in the literature, as described in Section 1, where each of these variables has been used by one or multiple other works to make a split in disk and elliptical galaxies. For consistency, many variables are remeasured from the particle data (in TNG300 and EAGLE) using a code developed for this reason, even if similar variables are available in simulation sub-halo catalogues. The variables needed to reproduce the results in this paper are publicly available for TNG300555https://doi.org/10.5281/zenodo.18740718, EAGLE [^6] https://doi.org/10.5281/zenodo.18740922 and Horizon-AGN [^7] https://doi.org/10.5281/zenodo.18740433.
The rotational energy divided by the kinetic energy of the stellar particles, $\kappa_{\mathrm{rot}}$, can be used as a proxy to determine whether a galaxy is rotation or dispersion dominated. This is often used as a definition for disk (rotation dominated) and elliptical (dispersion dominated) galaxies. $\kappa_{\mathrm{rot}}$ is measured in TNG300 and EAGLE and is defined as follows:
$\kappa_{rot,s}=\frac{\sum^{N_{s}}_{n=1}\left[\frac{1}{2}m_{n}(\hat{\boldsymbol{L}}_{s}\times\hat{\boldsymbol{r}}_{n})\cdot\boldsymbol{v}_{n}\right]}{\sum^{N_{s}}_{n=1}\left[\frac{1}{2}m_{n}\boldsymbol{v}_{n}\cdot\boldsymbol{v}_{n}\right]}$ (13)
@@
where the $^\hat{}$ -symbol denotes the direction of the vector and $\boldsymbol{v}_{n}$ is the velocity of each particle $n$ in galaxy $s$ relative to the centre of velocity of the galaxy, defined by:
$\boldsymbol{v}_{n}$ $=\boldsymbol{\dot{x}}_{n}-\boldsymbol{\dot{x}}_{s}$ (14)
$\boldsymbol{\dot{x}}_{s}$ $=\frac{1}{M_{s}}\sum^{N_{s}}_{n=1}m_{n}\boldsymbol{\dot{x}}_{n}$ (15)
@@
with stellar particle velocity, $\boldsymbol{\dot{x}}_{n}$ and mass-weighted velocity of the stellar particles for each galaxy, $\boldsymbol{\dot{x}}_{s}$. The angular momentum of the galaxy is given by:
$\boldsymbol{L}_{s}$ $=\sum^{N_{s}}_{n=1}m_{n}\boldsymbol{r}_{n}\times\boldsymbol{v}_{n}\ .$ (17)
@@
$|v/\sigma|$ is the rotational velocity divided by the velocity dispersion of a galaxy, and therefore based on the same idea of defining disks and ellipticals by kinematics as $\kappa_{\mathrm{rot}}$. It is measured following the procedure in Dubois et al. (2016). First, a cylindrical coordinate system with the angular momentum (Eq. 17) as $z$-axis and all coordinates ($R,\phi,z$) relative to the galaxy centre of mass and velocity ($\boldsymbol{r}_{n}$, $\boldsymbol{v}_{n}$) is adopted. Second, the mass-weighted mean of the $\phi$ component of the velocity is computed,
$\overline{v_{\phi,s}}=\frac{1}{M_{s}}\sum^{N_{s}}_{n=1}m_{n}v_{\phi,n}\ .$ (18)
@@
$\overline{v_{k,s}}$ $=\frac{1}{M_{s}}\sum^{N_{s}}_{n=1}m_{n}v_{k,n}$ (19)
$\overline{\Delta v_{k,s}}$ $=\frac{1}{M_{s}}\sum^{N_{s}}_{n=1}m_{i}(v_{k,n}-\overline{v_{k,s}})^{2}$ (20)
$\sigma$ $=\sqrt{\frac{\overline{\Delta v_{R,s}}+\overline{\Delta v_{\phi,s}}+\overline{\Delta v_{z,s}}}{3}}$ (21)
where $k$ is $R,\phi$ or $z$ and $n$ runs over the particles of galaxy $s$ as before. $|v/\sigma|$ is then defined as the absolute value of $\overline{v_{\phi,s}}/\sigma$.
The photometric colours used in this work are all from the available catalogues described in Nelson et al. (2018) for TNG300, Trayford et al. (2015) for EAGLE and Kaviraj et al. (2017) for Horizon-AGN, and measured to be SDSS-like, mimicking the colour filters of the Sloan Digital Sky Survey (SDSS), including dust extinction. Using colours to split a sample morphologically is one of the most widely used methods in simulations and observations.
Figure 1: Normalised distributions of $\mathrm{BTR}$ (top left), $|v/\sigma|$ (top right), $\kappa_{\mathrm{rot}}$ (bottom left) and rir-i (bottom right), which are used to split galaxies according to morphology, in TNG100 (light blue), TNG300 (dark blue), EAGLE (orange) and Horizon-AGN (green) at $z\approx 0$. Vertical lines have been added to indicate the split between disks and elliptical galaxies. The variables show a range of distributions with varying features across simulations. The colour bi-modality in TNG simulations is not present in EAGLE and Horizon-AGN. Most likely this is because they produce many low-mass red galaxies, due to limited resolution, that are not included in this mass cut.
The bulge-to-total ratio (BTR) is another variable based on the kinematic distinction between disks and ellipticals, by assigning stellar particles to the disk or bulge component of a galaxy based on their kinematic properties. This variable is not available for EAGLE and Horizon-AGN, and taken directly from the TNG300 catalogue (Genel et al., 2015), and therefore available for a limited number of galaxies, i.e. $M_{\star}>10^{9.3}\mathrm{M_{\odot}}/h$ at $z=0$ and $M_{\star}>10^{9.8}\mathrm{M_{\odot}}/h$ at higher redshifts. Therefore, $\mathrm{BTR}$ is included only in Section 3.1, and then examined separately in Appendix D. The $\mathrm{BTR}$ defined as 11- the fractional mass of stars in a galaxy with $\epsilon>0.7$, and measured as follows. First, the angular momentum of the stars within 10 times the stellar half-mass radius is computed. Second, the system’s z-axis is aligned with this angular momentum. Then, the stellar particles taken into account are sorted by their binding energy. For every stellar particle, the specific angular momentum, $J_{z}$ is calculated. $J(E)$ is noted, the maximum angular momentum of the stellar particles between 50 before and 50 after the particle in question (in the sorted list). Then, $\epsilon$ is computed for each particle by $\epsilon=J_{z}/J(E)$. Finally, the bulge to total ratio is computed as defined above.
@@ 3.1 Distributions
Figure 1 shows a comparison between the normalised distributions of the bulge-to-total ratio, $\mathrm{BTR}$ (top left panel), $|v/\sigma|$ (top right panel), $\kappa_{\mathrm{rot}}$ (bottom left panel) and the colour rir-i (bottom right panel), which are described in Section 2.4, for the simulations TNG100 (light blue), TNG300 (dark blue), EAGLE (orange) and Horizon-AGN (green) at $z=0$ ($z=0.0556$ for Horizon-AGN), when available. We use these variables in following sections to split the galaxy samples in disks and ellipticals, along the values added as horizontal dashed lines in the same colour.
Table 2: Number of galaxies in each sample for each simulation.
Redshift Simulation Positions Ellipticals Disks
0 TNG300 407180 20695 185880
Horizon-AGN 115208 6568 59112
EAGLE 22279 553 4972
1 TNG300 407568 18291 164660
Horizon-AGN 141546 6654 59887
EAGLE 22897 784 7062
The distributions of the variables that we use to perform the splits into disk and elliptical galaxy samples vary between simulations (Fig. 1 - 4). Therefore, we opt for the abundance matching method to define our samples, i.e. using a set elliptical fraction of $10\%$ of the total sample of galaxies with well-resolved shapes, as simulations tend to produce less large ellipticals than found in observations (Haslbauer et al., 2022). In line with the consistent stellar mass cut for the samples across redshift, we also fix the elliptical fraction, although we know observations have found both parameters to evolve with redshift (Mortlock et al., 2013). Probing the redshift evolution of a consistent sample of galaxies, e.g. using merger trees, is outside the scope of this work. The split value has been added as a vertical line for each simulation and variable in Figures 1 - 4. Table 2 contains the number of galaxies in each sample, where the total shape sample is equal to disks + ellipticals and the selection using the variables in Fig. 1 determines which galaxies make up each sample.
Figure 1 shows that the simulations agree on the ranges of the variable values, sometimes by definition, but not on how the variables are distributed. For TNG100 and TNG300, these differences can be ascribed to resolution differences that we explore in Appendix A, together with a discussion on the influence of boxsize and resolution on the results.
@@ 1 TNG300 407568 18291 164660
The differences between TNG, EAGLE and Horizon-AGN cannot be as easily explained as the differences between simulations in the same suite. Apart from varying resolutions and boxsizes, these simulations have different models and subgrid-physics recipies, as described in Section 2.1. Furthermore, where the kinematic parameters, $\kappa_{\mathrm{rot}}$ and $|v/\sigma|$, may be similar between TNG and EAGLE, the distributions of the colours have diverging features, of which only rir-i is shown in Figure 1. TNG displays a bi-modality in its colours, as found in observational surveys, e.g. GAMA (Taylor et al., 2015) and SDSS (Blanton et al., 2003), whereas this feature is not present in EAGLE or Horizon-AGN. Colour bi-modality has been found in Horizon-AGN (Kaviraj et al., 2017), albeit weaker than in observations. As there are many low-mass red galaxies in Horizon-AGN (Kaviraj et al., 2017), our mass cut is a likely cause of the uni-modality seen in Figure 1. EAGLE also contains too many low-mass red galaxies due to limited numerical resolution. For a detailed comparison between GAMA and EAGLE, see Trayford et al. (2015).